### v2

# Topic-based Watermarks for LLM-Generated Text

Abstract.
Recent advancements of large language models (LLMs) have resulted in indistinguishable text outputs comparable to human-generated text. Watermarking algorithms are potential tools that offer a way to differentiate between LLM- and human-generated text by embedding detectable signatures within LLM-generated output. However, current watermarking schemes lack robustness against known attacks against watermarking algorithms. In addition, they are impractical considering an LLM generates tens of thousands of text outputs per day and the watermarking algorithm needs to memorize each output it generates for the detection to work. In this work, focusing on the limitations of current watermarking schemes, we propose the concept of a “topic-based watermarking algorithm” for LLMs. The proposed algorithm determines how to generate tokens for the watermarked LLM output based on extracted topics of an input prompt or the output of a non-watermarked LLM. Inspired from previous work, we propose using a pair of lists (that are generated based on the specified extracted topic(s)) that specify certain tokens to be included or excluded while generating the watermarked output of the LLM. Using the proposed watermarking algorithm, we show the practicality of a watermark detection algorithm. Furthermore, we discuss a wide range of attacks that can emerge against watermarking algorithms for LLMs and the benefit of the proposed watermarking scheme for the feasibility of modeling a potential attacker considering its benefit vs. loss.
Large Language Models (LLMs), Natural Language Processing (NLP), Watermarking
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
The rapid expansion of artificial intelligence (AI) and natural language processing (NLP) in recent years has allowed for the development of large language models (LLMs) to become a prominent aspect in academia and industry (Inc., 2024). Models such as OpenAI’s ChatGPT (OpenAI, 2024), enables capabilities for generating human-like text by understanding user input and learning from extensive amounts of data (Ray, 2023; Zhao et al., 2023b). As more users are drawn to these systems, the amount of LLM-generated text over the internet increases, causing potential misinformation, copyright, and plagiarism issues (Rillig et al., 2023; Kalla and Kuraku, 2023). These concerns result in the need for tools that can differentiate LLM- and human-generated material.
Watermarking algorithms have potential to provide ways in which one can identify LLM-generated text by embedding an identifiable pattern in the text sequence (Srinivasan, 2024). Current watermarking schemes for LLMs involve adjusting the output token (i.e., units of text used in NLP analysis, such as words, phrases, or characters as elements for processing in NLP tasks) distribution strategically by separating tokens into “green” $G$ and “red” $R$ lists and enhancing the likelihood of $G$ listed tokens over $R$ listed tokens to be used in the LLM output (Kirchenbauer et al., 2023; Zhao et al., 2023a). The detection of a watermarked LLM- or human-generated text sequence is determined by the number of $R$ listed tokens in the text sequence. A human-generated output is expected to contain significantly more tokens in the $R$ list than in the $G$ list, the opposite is true for LLM-generated output containing more $G$ listed tokens compared to the number of $R$ listed tokens.
One major limitation of current watermarking algorithms is the lack of robustness against different attack models concerning text insertion, manipulation, substitution, and deletion. In such attacks, the goal of a potential attacker is to tamper with the LLM output to remove or distort the watermark, so the watermark detection mechanism cannot identify LLM-generated output with high confidence. In addition, efficiency and practicality limitations arise with the growth of LLMs and their usage. The creation of $G$ and $R$ lists for a watermark becomes impractical because the detection mechanism’s computational load requires it to exhaustively examine all seeds used to randomly partition the vocabulary of the LLM into a $G$ list the generated output is sampled from. With current model schemes, the model owner memorizes each query received in order to differentiate between an LLM- and human-generated text sequence by using the $G$ and $R$ pair of lists.
To resolve the aforementioned limitations of current watermarking algorithms, we propose the concept of a new model to enhance the robustness and efficiency by utilizing $G$ and $R$ lists in a topic-based manner. Using an input text sequence or the output from a non-watermarked LLM output, topics (e.g., sports, health, politics, technology, etc.) are extracted to generate a pair of lists for each topic. The designated lists for each topic are employed to detect whether a target text sequence was generated by an LLM or by a human. The practicality of the proposed scheme allows the LLM owner to disregard the infeasible task of iterating over all possible lists or seeds for the list pairs of an individual text input sequence and to generate only topic-based seeds for pairs of lists to watermark the text output sequence. On the other hand, potential attacks against watermarking algorithms mainly aim at tempering with the watermark in such a way that the detection algorithm fails to detect the watermark. Using a topic-based watermarking algorithm for LLMs also makes it easier to model the benefit vs. loss of such attacks. For instance, if an attacker, aiming to distort the watermark, modifies the LLM output in such a way that the topic of the output changes, and hence the detection algorithm fails with high probability, also significantly reduces the utility of the LLM output. Therefore, topic-based watermarking can be used to quantify the (i) success of the attacker in cheating the detection algorithm and (ii) decrease in the utility of the LLM output due to the attack.
The focus of this article is to discuss current watermarking algorithms, highlight their limitations, and provide the concept of a new watermarking scheme that mitigates such limitations. Specifically, we examine the functionality of contemporary watermarking and extendable robustness schemes, including an algorithmic and limitation standpoint for each scheme. We then propose an overview of our topic-based watermarking scheme for LLM-generated text, how our model can address current limitations, and limitations of the proposed model itself. Finally, we present future research and studies to empirically evaluate the proposed topic-based scheme.
2. Background
Introduced by Vaswani et al. (Vaswani et al., 2017), LLMs leverage a transform architecture which utilize attention mechanisms, allowing it to weigh relevant features (words or tokens) in a sentence, regardless of their positional distance in relation from one another (Galassi et al., 2019). Training an LLM involves feeding it extensive amounts of text data and adjusting internal parameters based on prediction errors the model makes (Shevchuk, 2023). The objective for training the LLM is focused around next-token predictions where the aim is to predict the next token in a sequence given the tokens which precede it.
The generation of tokens to output coherent text by an LLM is a sequential process, where each new token is produced based on the context provided by prior tokens. This process relies on a probabilistic model where for any given sequence of tokens, the LLM calculates the probability distribution over its vocabulary for the next token in the sequence. Different strategies the LLM employs to select the next token such as, selecting the token that has the highest probability (greedy or naive methods) (Dehqan, [n. d.]). The probabilities associated with each next token are derived from the LLM’s logits. Logits are the raw, unnormalized outputs of the LLM’s final layer (direct outputs from the LLM’s neural network computations before any normalization has been applied) which yields the probabilities for each token in the vocabulary. These probabilities reflect the model’s learned associations and its confidence in different continuations of the output text sequence.
Watermarking is a technique which involves embedding hidden markers or patterns within data to assert copyright, verify authenticity, and identity sources of data breaches. Research concerning the robustness of watermarking methodologies exists in a variety of fields. In prior work, Ji et al. (Ji et al., 2022) have focused on robust watermarking methods for genomic databases, ensuring accountability when sharing genomic data. Their research specifically addresses the mitigation of correlation attacks (an attack aiming to detect and distort the added watermark) against malicious database recipients by changing or watermarking database entries. While multiple watermarking schemes have demonstrated robustness against various threat models (Ji et al., 2022; Jiang et al., 2022; Ji et al., 2023), theses approaches have limited applicability within the context of LLMs. This limitation has facilitated the adoption of watermarking concepts to the field of generative AI, including areas such as NLP and LLMs. In the context of LLMs, watermarkings can be applied either to the data processed by the LLM or to the outputs they generate. These watermarking schemes are created to be invisible to human individuals, but identified through algorithmic detection mechanisms, enhancing the security and verification of generated text sequences.
3. Related Work
In this section, we explore the initial watermarking schemes developed for LLMs and examine enhancements aimed at increasing the robustness of these watermarking strategies.
3.1. Watermarking Algorithms for LLMs
To the best of our knowledge, the first watermarking scheme utilizes a probabilistic pair of “red” $R$ (tokens that cannot be used) and “green” $G$ (tokens that can be used) lists for LLM-generated text output. The work by Kirchenbauer et al. (Kirchenbauer et al., 2023), offers the generation of two types of $R$ lists, a soft $R$ list $R_{s}$ and a hard $R$ list $R_{h}$, along with their respective $G$ lists. Both types of $R$ and $G$ lists are used in the detection of human- versus LLM-generated text output, comparing the number of tokens from the $R$ and $G$ lists to the target text sequence (e.g., output of an LLM or human individual).
The generation of the $R_{h}$ and $R_{s}$ lists involve utilizing the initial input prompt and a sequence the first tokens $t-1$ that the LLM has already produced. The two $R$ lists are designed to either generate probabilities (which represent the likelihood for each token within the LLM’s vocabulary) or logits (which are discrete probabilities over the LLM’s vocabulary), depending on the specific $R$ list ($R_{h}$ or $R_{s}$) being used. By utilizing the last known token over the input sequence, a hash is computed to seed a random number generator which is then used to randomly partition the vocabulary the into $G$ and $R$ lists.
The difference between the $R_{s}$ and $R_{h}$ lists is because the $R_{s}$ list utilizes the addition of a constant “hardness” parameter to each of the $G$ list logits. The hardness parameter allows the sampling distribution for the next token in the generated sequence to strongly favor the $G$ list for high entropy tokens. High entropy tokens are a sequence of tokens that are good choices for the next token in the generated sequence, while low entropy tokens consist of one or few good choices of tokens to choose from as the next token in the generated output. The detection of the watermark for both $R$ lists are similar and it is done by checking each token in a target text sequence and examine for violations in the $R$ list.
There are limitations concerning the $R$ list generation of the watermarked LLM-generated output. To be able to detect the watermark, the $R$ lists need to be correctly reproduced specific to the generated output text. Limitations arise if the lists cannot be properly generated which may lead to incorrect conclusions on the analyzed target text (i.e., whether it was LLM- or human-generated). Overhead on maintaining the watermarking scheme also adds complexity to the LLM. It becomes infeasible to exhaustively iterate over all memorized lists or their specific hash functions and random number generators due to the number of users propriety LLMs have experienced in recent years. Our proposed model is hypothesized to detect watermarked LLM-generated text sequences efficiency and feasibly, reducing the complexity of the size of seeds for each pair of lists, leading to accurate classifications between LLM- and human-generated output.
3.2. Robust Watermarking
Limitations considering robustness of watermarking schemes has led to an influx of research for robust watermarking algorithms (Smith et al., 2023; Liu et al., 2024, 2023; Tang et al., 2023; Christ et al., 2023; Li et al., 2024; Zhao et al., 2023a). The first watermarking algorithm by Kirchenbauer et al. (Kirchenbauer et al., 2023), to extend its robustness, describes a robust private watermarking algorithm. The robustness extension involves employing the same generative process of their previously mentioned watermarking scheme, with the addition of a single secret key or multiple secret keys. The secret key(s) enhances the security and robustness on the watermark by generating unique tokens for the $R$ list utilizing a pseudorandom function. The additions of the key(s) further address the limitation of reproducibility concerns for the $R$ and $G$ lists.
Motivated due to the lack of robustness in existing watermarking algorithms, Liu et al. (Liu et al., 2024) propose a watermarking algorithm to enhance attack robustness on different semantically invariant perturbations (e.g., text modifications, text alterations). They achieve this by embedding watermark based on semantics of the entire prior text instead of a fixed number of prior tokens. Their scheme maintain robustness against attacks such as text paraphrasing (rewriting the watermarked text while preserving its original meaning) and synonym substitution (certain watermarked words are replaced with their respective synonyms). Christ et al. (Christ et al., 2023) propose an undetectable watermarking scheme utilizing cryptographic one-way functions, making it computationally infeasible to distinguish a watermarked output compared to other non-watermarked text. The watermarking scheme embeds a secret in the output text to which it does not affect the text quality generated by the LLM. The detection algorithm of the undetectable watermarking scheme identifies the output of the watermarked LLM without any information other than the output text sequence and the secret key.
While research in robustness of watermarking schemes for LLMs is quickly advancing, to our knowledge, there is little work concerning actual commercialized scenarios in which a usable watermarking detection scheme is feasible. Li et al. (Li et al., 2024), propose a watermarking scheme that may be practical in the commercial settings, introducing different backdoor data paradigms which integrate the watermark to the LLM during the tuning process. The motivation of the authors is to guarantee the copyright of a customized LLM from a business perspective comparable to prior watermarking schemes concerning the copyright of genomic databases (Ji et al., 2022). Li et al. focus on the copyright of an LLM in an internal company environment instead of differentiating between human- and LLM-generated text sequences. Prior research has also assumed black-box detection approaches where the internal workings of the model are undisclosed to the user. Limitations for black-box detection schemes revolve around data biases where the data collection of LLMs can introduce biases impacting performance and the ability to generalize across different context of the detection mechanism, confidence calibration affecting the overall evaluation metrics of the LLM, indicating the reliability and trustworthiness of the LLM, and the lack of adaptability to advance the learning of the LLM (Tang et al., 2023). Our proposed watermarking methodology differentiates from existing literature by enhancing practicality and robustness against attacks, specifically through our approach to store $G$ and $R$ list pairs tailored to topics of the input text sequence. Aligning with current studies, we utilize list pairs to elicit specificity and robustness in watermarking LLM-generated text, ensuring the watermark remains contextually relevant and minimally intrusive. This adaptation not only preserves text quality but also complicates watermark removal, increasing security against attacks and reducing false positives in watermark detection. Our proposed watermarking scheme further seeks to reduce computational demands and improve resistance toward known threat models, offering a practical solution for watermarking LLM-generated text.
4. Proposed Method
This section provides an overview of the proposed watermarking framework and other aspects of the model to be considered. First, in Section 4.1, we discuss different relevant attack models. In Section 4.2, we introduce the highlights of our proposed watermarking framework while comparing it to the work of Kirchenbauer et al. (Kirchenbauer et al., 2023). Then, in Section 4.3, we discuss our watermark detection scheme to differentiate between human- and LLM-generated output. Finally, in Section 4.4, we touch on the limitations of our proposed watermarking methodology.
4.1. Threat Model
Embedded watermarks in an LLM’s generated output are subject to various types of attacks summarized in the following. Note that all considered threat models are those where the motivation of an attacker is to distort or remove the watermark from the target text sequence to avoid detection. On the other hand, distorting the text output sequence will degrade the overall quality of text, and hence the topic of the corresponding text. To assess the impact of manipulations on the text quality, we propose to utilize the evaluation metric, perplexity, seen in prior research (Kirchenbauer et al., 2023) to measure how well a probability model predicts a sampled token. The perplexity metric is relevant for quantifying the degradation in text quality due to unexpected tokens as a result of efforts to evade watermark detection.
While we examine a range of real-life applications for watermarking, we specifically stick to the generation and submission of coursework in academia as an illustrative example. This scenario serves to illustrate the broader relevance and necessity of watermarking techniques, without confining the discussion solely to academia. In this context, the quality of manipulated text is determined by assessing its coherence, measured by the evaluation metric of perplexity. This evaluation approach also examines whether the text sequence accurately addresses and answers specific questions related to the coursework. This approach highlights the importance of both integrity and accuracy in an academic context. We assume that, in order to remain undetectable to the detection mechanism, a malicious user will manually change tokens while aiming to preserve text quality in relation to the specified topic(s). However, in the case of coursework submissions, preserving the quality extends beyond topic relevance to include precision and accuracy of the responses of the posed questions in the coursework submission. Modeling of this additional quality is yet to be realized, but addresses the importance of general and context-specific concerns in watermark detection.
Potential tools of an attacker include text insertion, substitution, and deletion in the generated output text sequence at the character, word, sentence, and multi-model (combination of characters, words, or sentences) level (Kirchenbauer et al., 2023; Goyal et al., 2023). Insertion is the addition or insertion of tokens into the text sequence. With the watermarked text encompassing significantly more $G$ listed tokens, the addition of new tokens may violate the $R$ list, increasing the count of $R$ listed tokens with respect to the $G$ listed tokens. However, the attacker will need to insert a significant amount of $R$ listed tokens to decrease detection power of the watermark to classify the text as human-generated. For instance, in the context of coursework submission, given a 1-page LLM output for an assignment, a student can add additional (human-generated) pages to decrease the detection power. However, this involves a trade-off: if the student is able to write one or more pages worth of content manually to remove the watermark, there would be no need to use LLM-generated text at all. The scenario of insertion attacks is realistic, but the fact that an attacker would need to manually generate the same or more content than the LLM-generated output is unrealistic, hence we expect a reasonable amount of insertion (i.e., manually generated tokens) as a part of the insertion attack. Substitution of tokens, involve swapping tokens in the generated output text. For example, an attacker may substitute tokens in the text sequence with their respective synonym. The motivation of the attacker in an instance of substitution is to remove $G$ listed tokens and add $R$ listed tokens to degrade or remove the watermark. Deletion involves deleting tokens from a generated output sequence. If a significant amount of $G$ listed tokens are deleted, the amount of $R$ tokens would increase however, if only tokens are deleted from a text sequence, there exists a potential trade-off between the text quality and detection accuracy. Similar to insertion, we also expect a reasonable level of deletion in a rational attack scenario. Token insertion, manipulation, and deletion are used in tandem to distort or remove the watermark of an LLM-generated output. An attacker can use a combination of these aforementioned tools to generate its attack against the watermarking scheme. In the rest of this section, we discuss considered threat models an attacker can perform against the LLM in our evaluation.
4.1.1. Baseline Attack
Baseline attack consists of the insertion, substitution, and deletion of text for a given output sequence. The attacker selects a single or a combination of techniques (insertion, substitution, and deletion) with the objective to diminish detection accuracy. In doing so, this approach may lead to a compromise in the text quality of the manipulated sequence.
4.1.2. Paraphrasing Attack
A paraphrasing attack is a category of a baseline substitution attack. As previously stated, execution of this attack may be manual by an individual or by rephrasing the output via an LLM. The proposed watermarking scheme is robust against varying scenarios of an attacker paraphrasing the output from the watermarked LLM. We consider separate paraphrasing attack scenarios.
  • Paraphrasing using the original LLM: The attacker generates a watermarked text sequence by feeding the LLM an input prompt. The attacker does not manually alter any tokens in the output. Instead, the attacker subsequently supplies the watermarked output into the identical LLM which originally produced it, along with a paraphrase command (i.e., “rephrase this text prompt”). The proposed watermarking scheme is robust against paraphrasing using the original LLM due to the nature of topics that form the basis of the watermarking process. For instance, if the attacker’s initial prompt leads to an output about “sports” (in terms of its topic), any paraphrased output would maintain the same topic, ensuring the watermark (on the paraphrased output) can still be generated and detected in the same way.
  • Manual Paraphrasing: This time, the paraphrasing is conducted manually by the attacker without the aid of external tools. This manual paraphrasing approach is considered for shorter sequences of text, as longer sequences would necessitate more extensive intervention from the attacker. The proposed topic-based watermarking scheme is particularly effective against manual paraphrasing, especially for lengthier text sequences, which would require significant modification to mislead the detection mechanism into incorrectly classifying the LLM-generated text for human-generated text. The robustness of the proposed watermarking scheme is further enhanced by the potential deterioration in text quality and the possible change in the overall topics of the paraphrased text sequence.
4.1.3. Tokenization Attack
A tokenization attack is classified as a form of insertion, where a token is modified into multiple subsequent tokens. For example, given the token “sports” within a $G$ list. An attacker can add characters such as ‘_’ or ‘*’, creating additional tokens categorized as $R$ listed tokens by the detection scheme. The original $G$ listed token, “sports”, is manipulated into “s_p_o_r_t_s”, creating multiple new $R$ listed tokens in the text sequence. The robustness of the proposed topic-based watermarking scheme concerns the degradation in text quality due to the intentional addition of noise (unwanted characters) within tokens in the $G$ list.
4.1.4. Discrete Alteration Attack
An attacker’s motivation behind this attack is to induce alterations in tokens, subsequently introducing misspelling and grammar errors. These manipulations can be executed through insertion or deletion of single or multiple characters, or even entire stings of characters within the specified output sequence, with the aim of diminishing the effectiveness of the watermark in detection. We consider an attacker which seeks to undermine the integrity of a watermarked text sequence without completely altering the text. By inserting or deleting characters within certain tokens, the attacker can modify the text to introduce errors, while minor, can impact the detectability of the watermark. For example, altering the word “watermarked” to “wat3rmark3d” through character substitution introducing misspelling by complicating the process of watermark recognition. The proposed topic-based watermarking addresses current watermarking algorithm limitations through the implementation of the generalized topic scheme. Even in the presence of minor spelling or grammatical errors, the overarching topic of the output remains unaffected. This ensures that the watermark’s integrity is maintained, showcasing the robustness of the proposed approach against textual alterations.
4.1.5. Collusion Attack
A collusion attack begins with an original text sequence that has been produced by an LLM, which incorporates a watermarking scheme based on topic relevance. Ideally, we assume that all existing LLMs have a topic-based watermarking scheme implemented for output text generation. [^1] Without this assumption, all watermarking schemes are vulnerable with the help of an LLM that does not add a watermark to its output. The attacker then solicits outputs to the same prompt from several LLMs, aiming to blend these multiple responses int a single output through insertion, manipulation, or deletion. This process might modify the topics extracted from the text, depending on the extraction methods and the defined list pairs of general and specific topics ($G$ and $R$). Should any of the LLMs employ its detection mechanism for classification purposes, it would label the mixed outputs as “LLM-generated”, effectively countering collusion attacks by recognizing the manipulation. However, the involvement of multiple LLMs in this process disrupts the correlation between the $G$ and $R$ list pairs, rendering the detection of the text sequence challenging and potentially leading to incorrect classifications. While the use of multiple LLMs in text manipulation attacks presents a complex challenge for the detection mechanism due to the potential alteration of extracted topics and the breakdown of general and specific topic correlations, employing detection mechanisms can help identify LLM-generated content. Nonetheless, the effectiveness of these mechanisms might be compromised as the number of intervening LLMs increases, highlighting the need for advanced strategies in watermarking and detection to mitigate such threats.
4.2. Proposed Watermarking Scheme
As discussed in Section 2, LLMs generate the next token in the output text sequence from the previously generated token, specifically the work by (Kirchenbauer et al., 2023), the tokens are taken from the specified $G$ list. Our proposed watermarking scheme is inspired by this process, however our methodology differs with the utilization of extracted topics from a non-watermarked LLM-generated output (or the input prompt) and the generation of $G$ and $R$ list pairs based on the extracted topic so that the detection algorithm does not need to exhaustively try all previously generated $G$ and $R$ list pairs.
We propose the concept of topic-based $G$ and $R$ list pairs combined with a topic-based watermarking mechanism (TBWM). The proposed TBWM framework consists of two different LLM components, a watermarked and non-watermarked LLM. The TBWM extracts the topic from the input prompt or a non-watermarked LLM-generated output based on the input text sequence. Figure 1 describes the proposed topic-based watermarking scheme.
Figure 1. Topic-based watermarking mechanism (TBWM). The input prompt is passed to non-watermarked LLM to extract the topic(s) of interest. Utilizing the topic(s), token weights are adjusted for the watermarked LLM with the original non-watermarked output.
4.2.1. Non-Watermarked LLM Topic Extraction
The proposed method for topic extraction employs an non-watermarked LLM output. This process initiates with a user prompting the LLM with an input prompt. Subsequently, this input sequence is input to the LLM which produces a non-watermarked output text sequence. This output serves as the basis for extracting topics relevant to the input prompt provided by the user. Furthermore, the non-watermarked output is then fed into the LLM in combination with a command to rephrase the text, such as “rephrase this text.” Both the non-watermarked output and the adjustments in token weight, derived from the extracted topics, are employed within the watermarked LLM to generate a watermark on the output text sequence. A secret key is used at this stage to influence the creation of the $G$ and $R$ list pairs. The lists are generated based on a pseudorandom function that incorporates the secret key, ensuring the watermarking scheme remains unpredictable and more robust against unauthorized detection. The utilization of the non-watermarked output over the input text sequence is to extract accurate topics when (i) a small input prompt occurs, consisting of a couple words to a sentence in length, and (ii) when misspellings and other grammar issues are present. The non-watermarked output of the LLM allows for a sufficient amount of quality text to be generated in order to extract an accurate topic, identifying the topic of interest of the corresponding user. Furthermore, for a single input text sequence, the watermarking scheme can extract multiple topics. For instance, in an input asking for medical injuries in different types of sports, the non-watermarked output could summarize the generalized topics from this specific output as “sports” and “medical injuries.” However, as to not degrade model performance or topic extraction quality, given a large non-watermarked output (e.g., paragraph(s)), the topic extraction may iterate over all sentences in the text sequence to analyze for relevant topics and skip those that are not strongly related to specified topics. The extracted topic or combination of topics are then used for seed and $G$ and $R$ list pair generation based on the number of extracted topics. The $G$ and $R$ list pairs are utilized to tune the token weights for the watermarked LLM along with the prior non-watermarked LLM-generated output that was used for topic extraction. The watermarked LLM utilizes the non-watermarked output text sequence in combination with a rephrase command to correctly sample from the generated $G$ list per topic. Similar to the watermarking algorithm proposed by Kirchenbauer et al. (Kirchenbauer et al., 2023), we sample tokens from the $G$ list to generate an output sequence from the original input prompt.
Via a similar process, we consider an alternative topic extraction method, directly from the user’s input text sequence. The extraction of topics from the input text is feasible when the text is free from spelling and grammatical errors and contains enough text to accurately extract topics. Simultaneously, the input text is processed through a non-watermarked LLM, resulting in a non-watermarked output based on the said input. The non-watermarked output serves as the input for the watermarked LLM along with the output and the adjustments in token weight, derived from the topics extracted from the input text sequence, are utilized within the watermarked LLM to produce a watermarked output text sequence. The approach where topics are employed to generate the $G$ and $R$ list pairs remains constant across both the non-watermarked topic extraction and the input prompt topic extraction methods.
The utilization of the global topic-based $G$ and $R$ lists allow for the watermarking of an input text sequence to be computationally sufficient by generating certain lists based on the topics allotted by the model. This generalization of $G$ and $R$ lists significantly decreases the overall number of list pairs needed for the detection of the watermark.
4.3. Watermark Detection Mechanism
Existing watermarking algorithms that rely on the $G$ and $R$ list pair methodology are based on the occurrence of $G$ and $R$ listed words within the target text (LLM output text sequence) described in Section 3.1. This detection method focuses on the frequency of $R$ list violations within the target text. Text produced by human individuals typically exhibit a higher usage of $R$ listed tokens, resulting in fewer instances of $G$ listed tokens in the target text. Conversely, watermarked LLM-generated text includes significantly more $G$ listed tokens over $R$ listed tokens. However, it is important to mention the infeasibility of this approach due to the detection method’s need to exhaustively try different list pairs or seeds of list pairs, complicating its practical application. Kirchenbauer et al. (Kirchenbauer et al., 2023) demonstrated the low likelihood of an individual being able to generate a text sequence, aligning with $G$ listed tokens similar to that of a watermarked LLM-generated output.
The proposed topic-based detection framework shown in Figure 2, mirrors the approach of leveraging $G$ and $R$ token list pairs. However, it introduces a novel mechanism classifying text as either human- or LLM-generated by extracting and analyzing the topics present in the non-watermarked LLM output sequence. The overall process involves the creation of specific $G$ and $R$ list pairs for each identified topic allowing the detection mechanism to iterate through all list pairs per topic to perform classification. Specifically, upon receiving a non-watermarked LLM output sequence, the detection framework identifies one or multiple topics present within the input text, utilizing topic modeling techniques (e.g., Latent Dirichlet allocation, non-negative factorization) (Nath, 2023). For each of the topics identified, the framework generates the pair of $G$ and $R$ lists, quantifying the occurrences of token from the pair of lists within the text. Given the distributions of $G$ and $R$ tokens across multiple different topics within the input text sequence, we employ hypothesis testing to determine the target text’s classification. The specific statistical tests chosen depends on the distribution characteristics of the token occurrence, aimed at evaluating the confidence level, whether the target text is human- or LLM-generated. Kirchenbauer et al. (Kirchenbauer et al., 2023) utilizes a “one portion z-test” involving the $G$ list of tokens with the expected value of tokens $T$ as $T/2$ and the variance of $T/4$. The overall z-test is denoted as:
(1) $\begin{split}z=2(G-T/2)/\sqrt{T}\end{split}$
which we will evaluate in our proposed detection scheme. The topic-based approach allows for a generalized analysis of the target text, which we hypothesize will accommodate the multiple topic text sequences and provide a workflow incorporating statistical tests to ensure the accuracy of the watermarking scheme.
Figure 2. Detection mechanism of topic-based watermarking scheme. Token list pairs, generated from the extracted topics of the non-watermarked LLM output, are compared to the output text sequence of the watermarked LLM to determine the classification of the target text as human- or LLM-generated.
4.4. Limitations
The proposed watermarking scheme is designed with computational feasibility for enabling the generation of $G$ and $R$ list pairs. The practically of the model implementation will depend on the number of categories allotted for the TBWM. The anticipated trade-off to be maintained concerns computational feasibility and the text quality of the LLM. For instance, should the input prompt be categorized under “sports” without a corresponding comprehensive list on the specific topic, the LLM’s output quality may be compromised. As granularity of the global list pairs increases, so too does text quality at the expense of increased computational demand and diminished model feasibility.
Although not the focus of this article, it is important to acknowledge the potential for spoofing attacks, a significant threat model that compromises the privacy and security of watermarked LLMs. These attacks deceive either the LLM or its users into believing that a given text output sequence originates from a different source (different LLM). Previous research has showed the susceptibility of LLMs to such attacks, with studies by Sadasivan et al. (Sadasivan et al., 2023), investigating the capability of attackers to generate derogatory outputs from LLMs by deducing text signatures without access to the detection mechanisms. Conversely, Pang et al. (Pang et al., 2024) concentrated on the watermarking dimensions, fabricating outputs that erroneously appear watermarked, both with the underlying motive of tarnishing the LLM’s or its developers’ reputation. Furthermore, the robustness of the proposed topic-based watermarking scheme against certain attacks represents notable limitations (Kirchenbauer et al., 2023; Pang et al., 2024). Specifically, the scheme is vulnerable to attacks aimed at removing the watermark through alterations in the output sequence, such as the addition, modification, or deletion of tokens. Despite these constraints, the topic-based approach, leveraging generalized $G$ and $R$ lists pairs, could potentially offer enhanced robustness through adjustment of weights within the global topic lists. The approach contrasts with generating a unique pair of lists for each output, potentially providing a layer of defense against both direct alterations and the threat of spoofing attacks by making the model less predictable and harder to manipulate. The resilience of this scheme to manual modifications in the output, as well as its effectiveness in defense against spoofing attacks that exploit the model’s sensitivity to topic predictions, will be subjects for future exploration, aiming to develop a watermarking solution that balances the demands of security, practicality, and text quality.
Furthermore, this work will understand the trade-off that emerges from refining granularity of generalized comprehensive list pairs with the topic-based watermarking scheme. The refinement of this trade-off, while beneficial for enhancing the specificity of the text output sequence, unintentionally introduces increased sensitivity to the accuracy of topic prediction. Sensitivity could potentially lower the threshold for adversaries to exploit the LLM by manipulating the model to bypass the watermarking scheme or produce false output sequences. The limitation falls in achieving the optimal balance to maintain the integrity of the watermarking framework but also adhere to the possible manipulation which could compromise the model’s feasibility. The evaluation for this trade-off will involve quantitative assessments of computational load and text quality and enhanced robustness against mentioned attacks (paraphrasing, tokenization, discrete alterations, etc.). Refining the global list pairs and the limitations of the proposed model, further motivate for a scheme consisting of ideal granularity and the ability to adapt to evolving threat landscapes, preserving the quality of output text sequences.
5. Conclusion
In this work, we propose a computationally feasible topic-based watermarking scheme that generates pairs of “green” and “red” lists based on an input text prompt topic of an LLM. We provide background knowledge of current watermarking algorithms and discuss their limitations being the motivation of the proposed framework with the lack of robustness against known attacks and the impractically of the memorization of text outputs. For future work, will involve the implementation and evaluation of the proposed watermarking scheme with the focus of improving constraints in terms of robustness and computation. Furthermore, there is potential for the expansion of our proposed watermarking scheme beyond a topic-based approach, such as generalizing other aspects of how lists are generated to detect text sequences.